A model airplane is to be an exact replica of the original, but at 1/50 the size. If the actual length of the airplane is 230 feet, how many feet long should the model be?
Answer by stanbon(75887)

A model airplane is to be an exact replica of the original, but at 1/50 the size. If the actual length of the airplane is 230 feet, how many feet long should the model be?
-----------------------------
x/230 = 1/50
x = (1/50)*230
x = 4.6 ft.
---------
Cheers,
Stan H.